OpenStudy (anonymous):

Can someone help me in here with this: |5 – 2x|= 9

OpenStudy (anonymous):

5 – 2x = 9 or -9

OpenStudy (mertsj):

5-2x=9 or 5-2x=-9 Solve those two equations.